OpenStudy (anonymous):

A restaurant charges $11 for a buffet plus $2 for each drink. Another restaurant charges $14 for a buffet plus $1 for each drink. The system of equations is graphed below. Part 1: What is the solution? Part 2: Show your work to verify the solution for both equations

ganeshie8 (ganeshie8):

where is the graph ?

ganeshie8 (ganeshie8):

@pinappleloveraries

OpenStudy (nurali):

11 + 2d = 14 + 1d 2d-1d=14-11 d=3 If two drinks are ordered, at both restaurants, the total will come to $17. 11 + 2(3) = 14 + 1(3) 11 + 6 = 14 + 3 17 = 17